OpenStudy (anonymous):

Valentina invested $6500 in a savings account with a yearly interest rate of 4% for 7 years. How much simple interest did she earn?

jimthompson5910 (jim_thompson5910):

I = P*rt I = 6500*0.04*7 ... plug in P = 6500, r = 0.04, t = 7 I = ??

OpenStudy (anonymous):

\[0.04\times6500\times7\] = 1820